OpenStudy (anonymous):

Simplify. (5 + i)2 A) 24 B) 24 + 10i C) 25 + 10i D) 26 + 10i

OpenStudy (anonymous):

\[(5 + i)^2=(5+i)(5+i)=(25-1)+(5+5)i=24+10i\]